### v2

# Bridging the Visual Gap: Fine-Tuning Multimodal Models with Knowledge-Adapted Captions

Abstract
Recent research increasingly focuses on training vision-language models (VLMs) with long, detailed image captions. However, small-scale VLMs often struggle to balance the richness of these captions with the risk of hallucinating content during fine-tuning. In this paper, we explore how well VLMs adapt to such captions. To quantify caption quality, we propose Decomposed NLI (DNLI), an evaluation framework that breaks down generated captions into individual propositions, assessing each in isolation. This fine-grained analysis reveals a critical balance between capturing descriptive details and preventing hallucinations. Our findings show that simply reducing caption complexity or employing standard data curation techniques does not effectively resolve this issue. To tackle this challenge, we introduce Knowledge Adapted (KnowAda) fine-tuning, a data-centric approach that automatically adapts training data with the model’s existing knowledge and visual understanding. KnowAda minimizes hallucinations while preserving high descriptiveness. We validate this approach across several small-scale VLMs (up to 7B parameters) and dense caption datasets, demonstrating that KnowAda effectively balances hallucination reduction and descriptiveness. Our results show that KnowAda outperforms various baselines in both automatic metrics and human evaluations. We will release our code and models.
Bridging the Visual Gap: Fine-Tuning Multimodal Models
with Knowledge-Adapted Captions
[…]
Figure 1: KnowAda identifies knowledge gaps of a VLM and adapts the dense caption accordingly. The KnowAda dense captions are better suited for downstream fine-tuning of the VLM.
Figure 2: Our proposed KnowAda pipeline. We first probe the knowledge of the VLM, identifying the known and unknown parts of the image description, by generating questions about the visual content of the image mentioned in the caption. Then, KnowAda identifies the knowledge gaps by judging the VLM answers to these questions. Finally, KnowAda adapt the description to match these gaps (e.g., removing the number of limousines mentioned in the caption, which relates to a question the model failed to answer).
1 Introduction
Fine-tuning pretrained multimodal models for generating dense image captions is common in both research and practical applications, such as assisting visually impaired individuals. Recent work has focused on creating high-quality, descriptive captions through human annotations Onoe et al. (2024); Garg et al. (2024); Deitke et al. (2024) and synthetic generations from models like GPT-4 Chen et al. (2023a, 2024a) and Gemini Singla et al. (2024), including extensions that integrate expert models for enhanced detail Li et al. (2024). These datasets enable the fine-tuning of models to create detailed descriptions in specific styles, distinguishing them from the zero-shot capabilities of pretrained models. However, smaller multimodal models (e.g., up to 7 billion parameters), which are essential for real-time applications, frequently face challenges in capturing fine-grained visual details during fine-tuning, resulting in hallucinations.
Consider a fine-tuning dataset for dense captioning, like the one in Figure 1 from DOCCI Onoe et al. (2024), alongside a pretrained vision-language model trained on tasks like captioning, VQA, and OCR. If, for instance, the model has only encountered low-resolution images during pretraining, it may struggle to identify fine details, such as the drawing on the purple van or the hotel name in the background, which require higher resolution. This issue extends beyond resolution to other visual challenges in modern VLMs Tong et al. (2024); Wu et al. (2024); Zhang et al. (2024a). We hypothesize that fine-tuning the model on overly complex captions may increase hallucinations, as the model is compelled to predict details it cannot accurately perceive or understand.
Recent works on large language models (LLMs) have shown that fine-tuning primarily adapts pre-existing factual knowledge for specific tasks, with the majority of this knowledge being encoded during the pretraining phase Geva et al. (2020); Meng et al. (2022); Roberts et al. (2020). Rather than acquiring new information, fine-tuning typically activates and refines pretrained knowledge, which remains largely stable throughout the process Zhou et al. (2024b). Furthermore, Gekhman et al. (2024) demonstrated that fine-tuning on content not grounded in a model’s pre-existing factual-knowledge can lead to an increase in hallucinations. Similarly, Yu et al. (2024) showed that attempting to distill GPT4V into a smaller, less capable VLM significantly increases hallucinations. Motivated by these findings, we hypothesize that one possible cause of increased hallucinations is the excessive visual complexity of captions relative to the model’s pretrained capabilities. We propose a method to mitigate this effect.
To better adapt pretrained models to dense caption datasets, we introduce KnowAda, a model-specific adaptation technique that simplifies complex details in dense captions. The KnowAda pipeline, shown in Figure 2, automatically identifies visual knowledge gaps between a pretrained VLM and an image-caption pair by generating questions related to the image’s visual content. It then modifies the captions to align with the model’s visual knowledge and capabilities, producing adapted captions that enhance fine-grained control while balancing a low hallucination rate with high descriptiveness.
Evaluating dense image captions requires attention to two critical factors: descriptiveness, which captures the image’s details, and hallucination rate, which measures factual accuracy. Traditional metrics, such as those based on similarity to reference captions Papineni et al. (2002); Banerjee and Lavie (2005); Zhang et al. (2019); Reimers (2019) or CLIP-based alignment Sarto et al. (2023); Radford et al. (2021), often fall short when applied to long captions. These approaches penalize valid variations in phrasing and fail to distinguish between factual accuracy and token overlap or semantic similarity, rendering them inadequate for identifying hallucinations in detailed captions. Existing hallucination metrics are similarly limited, focusing primarily on short captions or object-level errors Rohrbach et al. (2018); Li et al. (2023c); Ben-Kish et al. (2023), or relying on other VLMs Jing et al. (2023); Liu et al. (2023), which themselves can hallucinate during the verification process. To address these limitations, we propose Decomposed NLI (DNLI), a novel evaluation framework that breaks captions into propositions and assesses their entailment with the detailed ground truth description. DNLI offers a more reliable measure of both descriptiveness and accuracy, demonstrating strong alignment with human judgments.
Our results show that training with KnowAda captions offers a favorable balance between descriptiveness and fidelity when fine-tuning LLaVA-7B, outperforming other data-centric baselines. To demonstrate the consistency of KnowAda across different models and datasets, we fine-tuned several multimodal models, ranging from 2 billion to 7 billion parameters, on two different dense captioning datasets. Across all models, KnowAda consistently reduced hallucinations compared to training with the original captions, as confirmed by both automatic and human evaluations.
Stated explicitly, our contributions are: (I) We show that small-to-medium-scale VLMs underperform when fine-tuned directly on dense caption datasets, exhibiting increased hallucinations and reduced descriptive accuracy; (II) To address this we propose KnowAda, a model-dependent augmentation method that fills knowledge gaps in captions, reducing hallucinations while preserving high descriptive accuracy; (III) We demonstrate the effectiveness of KnowAda through extensive experiments, supported by both quantitative metrics and qualitative human evaluations; (IV) We introduce DNLI, a novel evaluation framework for dense captions that offers a more fine-grained analysis and shows strong correlation with human annotations.
Figure 3: Dense captioning descriptiveness precision-recall results for LLaVA-7B fine-tuned with DOCCI captions, adapted using different methods. “Trimmed” refers to naive removal of sentences, while “Gemini” involves prompting Gemini to simplify the caption by removing difficult details of varying degrees. KnowAda consistently achieves better precision-recall balance. Original captions corresponds to KnowAda with a threshold of $T=100\%$, where no information is classified as unknown.
Figure 4: DNLI Evaluation. Given a generated description by a VLM, we decompose it to atomic propositions. Then, we classify each proposition to either entailed, contradicted or neutral, conditioned on the ground-truth description. Finally, we calculate the descriptiveness and contradiction based on the number of entailed and contradicted propositions.
Contradiction $\downarrow$ Descriptiveness $\uparrow$
Model FT Captions Precision Recall Precision Recall # Words
Auto Human Auto Human Auto Human Auto Human
PaliGemma Synthetic 38.9 19 30 15.5 47.9 81 39.2 67.8 72
PaliGemma Synthetic KA 32.4 18.4 20.8 13 55.2 81.6 36.4 61.2 54
TinyLLaVA Synthetic 38.1 52.1 49.1 45 35.1 47.9 26.3 40.8 71
TinyLLaVA Synthetic KA 22.9 34 40.2 22.2 48.1 66 25.4 40.9 51
LLaVA-7B Synthetic 39.1 19.3 39.1 16.2 47.3 80.7 39.7 65.3 79
LLaVA-7B Synthetic KA 31 15.8 31 9.3 58.4 84.2 34.5 47.4 54
PaliGemma Human 41.6 20.4 24.6 14.9 46.7 79.6 24.6 43.2 62
PaliGemma Human KA 38.3 18.3 22.2 13.6 49.4 81.7 28.7 38.9 65
TinyLLaVA Human 53 51.8 39.5 38.8 31.9 51.8 22.4 31.4 100
TinyLLaVA Human KA 42.6 34.5 19.6 12.5 46.9 65.5 19.1 22.5 53
LLaVA-7B Human 47.2 33.4 39.7 33.2 39.4 66.6 33.7 48.1 109
LLaVA-7B Human KA 33.7 17.1 16.7 11.2 56.9 82.9 25.8 31.8 55
Table 1: Dense captioning results over the test sets of DOCCI when fine-tuning on original human-annotated captions, synthetic captions, and KnowAda-adapted captions (denoted as KA) with a threshold of 20%. “Automatic (Auto)” refers to model-based NLI evaluation, while “Human” refers to evaluations based on human labeling.
2 KnowAda
We begin by introducing our caption adaptation method for dense captioning datasets. KnowAda comprises three stages, as shown in Figure 2. First, we use an LLM to generate visual questions from each dense caption. Next, these questions are employed to probe the VLM’s pretrained visual knowledge and identify parts of the image caption the model struggles with. Finally, an LLM adapts the image descriptions by editing out the unknown parts. Below, we elaborate on each step of the pipeline.
2.1 VLM Knowledge Probing
In order to detect the visual attributes of the image that are unknown to the VLM but are mentioned in the image description, we probe the VLM knowledge. It is done by generating visual questions which can be answered by the image description, then letting the VLM answer these questions, and finally measuring if the responses are correct. We provide further details regarding each stage below.
Finding the unknown questions.  Given a dataset $D$ containing image descriptions, we aim to find all the parts of the description that are unknown to the VLM. Following prior work on uncertainty estimation Cheng et al. (2024); Gekhman et al. (2024), the steps are as follows:
1. For each image caption $d\in D$, generate $n$ questions $Q=\{q_{1},q_{2},\ldots,q_{n}\}$. Note that $n$ is caption dependent (e.g., more questions are generated for longer captions).
2. For each question $q_{i}\in Q$, sample $m$ answers. Let $A_{i}=\{a_{i1},a_{i2},\ldots,a_{im}\}$ represent the set of $m$ answers for question $q_{i}$.
3. Evaluate each answer $a_{ij}\in A_{i}$ to determine how difficult it is for the specific VLM. Let $C_{i}\subseteq A_{i}$ be the set of correct answers and $I_{i}\subseteq A_{i}$ be the set of incorrect answers for question $q_{i}$.
4. Calculate the difficulty of each question $q_{i}$ based on the ratio of incorrect to incorrect and correct answers. The difficulty score $df_{i}$ for question $q_{i}$ is given by:
$df_{i}=\frac{|I_{i}|}{|C_{i}|+|I_{i}|}$
Here, $|C_{i}|$ and $|I_{i}|$ are the number of correct and incorrect answers respectively. A higher value of $df_{i}$ indicates a more difficult question. For a threshold $T$, a question with $df_{i}>T$ is defined as an unknown question.
We assess the accuracy of the model’s responses by prompting a LLM to evaluate each generated answer in relation to the given question and the ground truth description.
For example, consider the two questions shown in Figure 2 (ii). If the model correctly answered ‘‘Monorail’’ for the question ‘‘What kind of vehicle is visible in the image?’’ in 6 out of 10 sampled instances, with 4 incorrect answers, the difficulty for this question is calculated as $df=\frac{4}{10}$ For the second question, ‘‘How many limousines are parked underneath the track?’’, the model correctly answered ‘‘3’’ only 4 out of 10 times, resulting in a difficulty of $df=\frac{6}{10}$. With a threshold of $T=50\%$, the first question is classified as known, while the second as unknown. However, if the threshold is raised to $T=70\%$, both questions would be considered unknown.
For example, consider the two visual questions shown in Figure 2 (ii) with the following prediction accuracies:
  • For the question ‘‘What kind of vehicle is visible in the image?’’, the model correctly answers ‘‘Monorail’’ in $6/10$ instances, resulting in $df=4/10=0.4$.
  • For the question ‘‘How many limousines are parked underneath the track?’’, the model correctly answers ‘‘three’’ in $4/10$ instances, yielding $df=6/10=0.6$.
With a difficulty threshold $T$, the classification of these questions changes:
  1. 1.
    For $T=50\%$:
    • The first question is Known (since $df=0.4<0.5$)
    • The first question is Unknown (since $df=0.6>0.5$)
  2. 2.
    For $T=70\%$:
    • The first question is Unknown (since $df=0.4<0.7$)
    • The second question is Unknown (since $df=0.6<0.7$)
Image description rewriting. After identifying the questions unknown to the model for each image description, we prompt a LLM, namely, Gemini with several in-context examples to remove or edit the parts of the caption that answer these unknown questions, while keeping the other details intact. For instance, in Figure 1, the questions about the number and type of cars parked in the parking lot is unknown to the small VLM. Therefore the corresponding questions are passed to Gemini, which edits ‘‘three limousines’’ to ‘‘vehicles’’.
Formally, for each image description $d_{i}$, we use all the corresponding questions classified as unknown to the VLM according to the threshold $T$, denoted as $Q{{}_{i}}=\{q_{i,1},q_{i,2},\ldots,q_{i,\bar{n}}\}$, to prompt the LLM to rewrite the description by removing information associated with $Q{{}_{i}}$. See Appendix A.1 for the prompts used at each stage.
2.2 KnowAda Data Analysis
We start by examining the characteristics of the dense captioning datasets, the impact of KnowAda on the adapted captions, and the types of questions utilized for visual probing within the KnowAda pipeline.
Datasets. We use two variations of dense caption datasets in our experiments: DOCCI Onoe et al. (2024), a human-annotated dataset rich in visual details, and DOCCI images paired with synthetic captions generated by Gemini-Pro-1.5, which are designed to be highly visually descriptive. These datasets differ in caption style and level of visual detail, allowing us to demonstrate the robustness of KnowAda across varying data distributions.
Dataset characteristics. Figure 6 illustrates the overlap of unknown questions across the different models. While there is a core set of unknown questions common to all three models, each model also has its own unique set of unknown questions.
Table 2 presents statistics for each dataset, including the average number of unknown questions per model and dataset, as well as the average word count in both the original and KnowAda captions. The data indicates that the average number of unknown questions and the average word count in the rewritten captions are relatively consistent across different models, where the human-authored DOCCI captions containing slightly more challenging questions compared to the synthetic captions.
Visual questions category distribution. To verify that KnowAda generates diverse types of visual questions, we use LLaMa-3-70B Dubey et al. (2024), quantized to 4 bits, to classify 12,422 questions generated for 1,000 images from the DOCCI test set into categories defined by SeedBench Li et al. (2023a). For the classification task, we provide one-few shot examples from each SeedBench category. The resulting distribution of questions is presented in Figure 5. As observed, the diversity in the question types suggests that KnowAda performs knowledge probing across a wide range of visual tasks.
3 Decomposed NLI Evaluation
Next, we introduce DNLI, a proposition-decomposition evaluation framework for dense-captioning. It evaluates the quality of a generated dense image caption using two criteria: descriptiveness, which measures accuracy and detail, and contradiction, which measures how much the caption contradicts the ground truth. Inspired by prior work on paragraph summarization evaluation Ernst et al. (2021); Zhang and Bansal (2021); Ernst et al. (2021) and retrieval Chen et al. (2023b), DNLI assesses descriptiveness and consistency through proposition extraction, as outlined below and shown in Figure 4. Additional qualitative examples are in the appendix.
Propositional decomposition. Given a generated image description, we use a Gemini to decompose it into a set of atomic propositions, capturing individual, verifiable claims. This enables fine-grained evaluation. To avoid duplicates, we instruct the model to include only unique propositions.
Natural Language Inference (NLI) analysis. We next evaluate the entailment of propositions generated from the image using both our automatic method, where Gemini is used to compute textual entailment, and through human annotators who directly assess visual entailment:
Automatic Textual Entailment:  Each atomic proposition is assessed with a Natural Language Inference (NLI) model. This model compares each proposition (the hypothesis) to a ground truth description of the image (the premise), determining whether the proposition is entailed, contradicted, or neutral with respect to the description. This evaluation forms the basis for our descriptiveness and contradiction metrics. See details and prompts in Appendix A.1.
Human-Based Visual Entailment Evaluation:  We employ Amazon Mechanical Turk to engage three independent annotators for each proposition, tasking them with directly assessing its entailment based on the corresponding image. To ensure quality, we administer a qualification test and select the top-performing annotators. Each proposition is evaluated by a distinct set of three annotators, and the final entailment label is determined by a majority vote. For consistency, we randomly sample 20 images per model, generating an average of 175 propositions per model. In total, 2,812 unique propositions were evaluated. Further details are provided in Appendix A.3.
Descriptiveness metric. We quantify the descriptiveness of the generated caption using two measures: recall and precision.
Descriptiveness Recall: The proportion of ground truth propositions that are entailed by the generated description:
Descriptiveness Recall=|Entailed||Ground $\text{Descriptiveness Recall}=\frac{|\text{Entailed}|}{|\text{Ground Truth}|},$
where the entailed propositions are those identified by the NLI model as described above.
Descriptiveness Precision: The proportion of entailed propositions relative to the total number of propositions in the generated description:
Descriptiveness $\text{Descriptiveness Precision}=\frac{|\text{Entailed}|}{|\text{Generated}|},$
where the generated propositions are those found in the model’s caption. The precision represents the likelihood that a given proposition extracted from the generated caption, would be entailed.
Contradiction metric. The contradiction precision and recall are calculated in a similar manner:
Contradiction Precision=|Contradicted||Ground $\text{Contradiction Precision}=\frac{|\text{Contradicted}|}{|\text{Ground Truth}|}$
Contradiction $\text{Contradiction Recall}=\frac{|\text{Contradicted}|}{|\text{Generated}|}$
Together, these metrics provide a comprehensive evaluation of the generated captions. Descriptiveness precision and recall assess the accuracy and coverage of the content, respectively. Contradiction precision indicates the likelihood of a false proposition in a caption, while contradiction recall measures the total number of contradictions.
Note that the neutral labels are discarded in the automatic evaluation, as they could represent either subjective propositions (e.g., "vibrant atmosphere") or visual claims not described in the original captions. These labels could be either entailed or contradicted, but it is unclear which.
Figure 5: Distribution of visual question categories generated from captions in the first stage of KnowAda, classified using SeedBench Li et al. (2023a) definitions.
Captions Model $\bar{C}_{o}$ $\bar{C}_{r}$ $\bar{Q}_{unk}$
Human PaliGemma 122 84.2 5.5
Human TinyLLaVA 122 80.9 6.3
Human LLaVA-7B 122 80.7 6.3
Synthetic PaliGemma 92 65.8 4.4
Synthetic TinyLLaVA 92 63.3 5.2
Synthetic LLaVA-7B 92 63.7 5.2
Table 2: Statistics of the original and KnowAda adapted captions with $T=20\%$ over the human annotated and synthetic version of DOCCI captions. $\bar{C}_{o}$ and $\bar{C}_{r}$ denotes the mean number of words in the original caption and caption after applying KnowAda, and $\bar{Q}_{unk}$ denotes the mean number of questions unknown to the model.
Figure 6: Overlap in $\bar{Q}_{unk}$ (average number of unknown questions per caption) across different models on the DOCCI training set for $T=20\%$. Each model has a unique set of unknown questions, with about 50% of these shared among all three models.
Method Contradiction $\downarrow$ Descriptiveness $\uparrow$ # Words
Precision Recall Precision Recall
KnowAda Random 34.7 17.2 57.9 23.9 55
KnowAda 33.7 16.7 58.9 25.8 55
Table 3: Design choices ablations.We ablate the effect of removing unknown information versus removing random information in the KnowAda pipeline. Removing unknown information improves performance across all metrics.
4 Experimental Settings
Our experiments concentrate on training dense captioning models. We evaluate the performance of models fine-tuned on captions curated through various methods against those fine-tuned on captions adapted using KnowAda.
We start by comparing KnowAda to data curation methods for multimodal datasets according to the DataComp challenge Gadre et al. (2024), specifically ICC Yanuka et al. (2024) and DFN Fang et al. (2023). We also compare our approach to the following baselines that might address the hallucination-descriptiveness trade-off:
  • Caption Trimming: A progressive method that removes varying numbers of sentences to simplify captions.
  • Gemini Simplification: An approach prompting Gemini to remove difficult details from captions to varying degrees.
For all methods in this experiment, we fine-tuned LLaVA-1.5-7B Liu et al. (2024) using Low Rank Adaptation Hu et al. (2021).
To ensure that KnowAda is robust across multiple models and datasets, we fix the threshold at 20% for classifying questions as unknown and fine-tune three models: PaliGemma Beyer et al. (2024), TinyLLaVA Zhou et al. (2024a), and LLaVA-1.5-7B Liu et al. (2024). We fine-tune on two variations of DOCCI: one using the original DOCCI captions, and another using synthetically generated captions created by Gemini, which were prompted to be visually descriptive. We evaluate the models using both an automatic NLI model and human annotators, as detailed in Section 3. In all experiments, we split the DOCCI test set into 1,000 samples for evaluation, while 4,000 samples are used for the reported test set.
5 Results
This section presents our experiments, highlighting the impact of KnowAda over competing baselines.
5.1 KnowAda Achieves Better Descriptiveness-Hallucination Trade-off
Figure 3 illustrates that fine-tuning on KnowAda captions consistently provides the best balance between high descriptiveness and low hallucination rates compared to competing baselines. While ICC and DFN offer slight improvements in precision, they do not facilitate further gains due to a lack of control over the trade-off. Although trimming and rephrasing captions allow for some control over precision and recall, they yield inferior results compared to KnowAda across all thresholds.
5.2 KnowAda Is Consistent Across Multiple Models and Datasets
Table 1 shows that fine-tuning on KnowAda captions consistently reduces the hallucination rate while maintaining high descriptiveness across different trained models. This superior performance is evident for both human-annotated and synthetically generated dense captions, as confirmed by our automatic and human evaluation pipelines.
Specifically, KnowAda significantly reduces the contradiction rate in terms of both precision and recall. However, while it improves precision in descriptiveness, it decreases recall. The results in Table 1 are obtained using a stringent threshold of $T=20\%$, prioritizing hallucination reduction. KnowAda allows increased recall by enabling the selection of a less stringent threshold, facilitating a trade-off between precision and recall, as illustrated in Figure 3. Future work may explore how to maintain KnowAda’s improvement in hallucinations while ensuring high recall in descriptiveness.
5.3 Textual Entailment is Highly Correlated to Human Annotated Visual Entailment
We computed the Phi correlation coefficient between the majority agreement labels (ground-truth human annotations) and the labels generated by the NLI model for each proposition. The results yielded a Phi coefficient of $\phi=0.73$ for the DOCCI original captions and $\phi=0.67$ for the synthetic captions. These strong positive correlations indicate a significant relationship between the human annotations and our proposed automatic evaluation. It suggests that our text-only NLI model effectively aligns with human judgments of visual entailment, demonstrating its reliability in distinguishing between contradictory and entailed propositions based solely on dense captions.
Moreover, the correlation difference between original and synthetic captions shows that DOCCI human-authored captions are generally more detailed and reliable compared to the synthetic captions, enhancing the effectiveness of the automatic entailment computation using them.
Note that the results from our human annotation process showed an average majority agreement of 82.6% for the “Contradicted” labels and 89.7% for the “Entailed” labels, indicating a high level of consensus among annotators.
5.4 Necessity of Removing Unknown Information
We perform an ablation study to assess the importance of removing VLM-unknown information by following the KnowAda procedure for a fixed $T=20\%$, but instead of removing information linked to unknown questions, we randomly remove information, a method referred to as KnowAda Random in Table 3. As expected, this results in worse performance across all metrics, demonstrating that unknown information is indeed a problematic factor in captions that impacts the hallucinations.
6 Conclusions
This work focuses on fine-tuning small-to-medium-scale vision-language models by aligning dense captions with the models’ existing knowledge and visual understanding. It aims to address the challenge of balancing rich descriptiveness with factual correctness in multimodal models, especially when models have limited capacity to process complex visual details in dense captions. By probing models with visual questions and adapting captions to exclude unknown information, we reduce hallucinations while maintaining high descriptive accuracy. Our results, supported by both human and automatic evaluations, suggest that this strategy can lower hallucinations compared to training on original captions or competing data curation baselines.
Furthermore, we introduce a proposition-based evaluation framework that provides fine-grained analysis of generated captions, offering deeper insights into the balance between descriptiveness and factuality. We believe that our findings contribute to improving fine-tuning methods in VLMs from a data-centric approach, particularly in resource-constrained environments where balancing descriptiveness and accuracy is important for real-world applications. We hope that this work will encourage further research into addressing the challenges of dense captions in VLMs.
7 Related Work
Our research is connected to the field of dense image captioning, particularly in addressing unknown information and enhancing fine-tuning datasets. Below, we provide an overview of previous works in each of these areas.
7.1 Dense Image Captioning
Traditional image captioning datasets like COCO Chen et al. (2015) often feature brief captions with limited visual detail. Recently, interest has grown in creating longer and more complex captions. For example, Onoe et al. 2024 and Garg et al., 2024 developed a dataset of 15K human-annotated, richly descriptive captions. Additionally, Chen et al., 2023a and Singla et al., 2024 employed GPT-4 and Gemini to generate detailed captions. We show that these datasets are more effective for fine-tuning when adapted to the specific model being trained.
7.2 Training On Unknown Information
Several works examine the relationship between unknown information and downstream model performance. Gekhman et al. (2024) show that fine-tuning LLMs on low-confidence examples encourages hallucinations, and suggest that fine-tuning should not introduce new knowledge, but only teach the model to make use of existing knowlegde. Zhang et al. (2024b) suggest optimizing for higher-confidence responses via Direct Preference Optimization (DPO), where the confidence of each response is estimated via self-evaluation. Piché et al. (2024) train a utility function that encourages predicting only high-certainty responses. The function is trained over synthetic data, which is collected iteratively, and is composed of fused high-confidence predicted facts. Additionally, Xu et al. (2024) introduce a refusal mechanism which encourages the model to reject questions that do not align with its existing knowledge. We demonstrate that training a model on questions it is fundamentally lacking the ability to answer, rather than solely on factual knowledge it lacks—particularly in the multimodal domain—results in similar performance degradation. Moreover, our method only rejects the unknown parts of a data sample, while maintaining the information that is useful for training, thus avoiding throwing out useful training samples which only include some unknown parts.
7.3 Data Curation in Fine-Tuning
Recent research has focused on refining the fine-tuning process for pretrained models, especially through instruction tuning. Zhou et al. (2024b) demonstrated that fine-tuning on a small, high-quality instruction dataset can yield superior results. Li et al. (2023b) explored training a language model on a subset of instruction data, using loss discrepancies to inform filtering strategies. However, this approach doesn’t extend to image captioning, where no instructions are available. Lin et al. (2024) proposed reducing noisy token influence by assigning lower weights, while Chen et al. (2024b) trained a network to filter out low-loss instruction samples. In contrast, our approach adapts specific parts of captions misaligned with the model’s capabilities, rather than filtering or reweighing the entire caption. Our strategy leverages visual knowledge and emphasizes sequences and conceptual coherence, moving beyond token-level adjustments.
8 Limitations
While KnowAda demonstrates notable effectiveness, several limitations must be considered.
Firstly, KnowAda is model-dependent, probing the knowledge of each model in isolation. This means that the data must be tailored for each model independently, leading to increased computational overhead.
Secondly, while we show that KnowAda achieves better performance in balancing the descriptiveness-hallucination trade-off, it does not fully resolve this issue. Reducing contradictions to the ground truth often decreases descriptiveness recall (but not precision).
Lastly, our analysis focused specifically on dense caption generation, which is a limited task. Expanding KnowAda to other tasks, such as visual question answering (VQA), could represent an interesting avenue for future research.
Ethics Statement
This work focuses on measuring and mitigating hallucinations in visual-language models (VLMs). As such, it is expected to increase the reliability of VLMs and the ability to measure their performance, which is important when using them in real-world systems. This is expected to have a positive impact on the use of VLMs in society. However, we recognize that the foundation models used in the KnowAda construction and evaluation pipeline could propagate biases. We anticipate further research into such biases before relying on our work beyond the research environment. The human annotation study performed in this work received the required IRB approval from our institution’s ethics committee.
Acknowledgments
This work was partially supported by Google and the TAU Center for Artificial Intelligence and Data Science. The authors thank Morris Alper and Nimrod Shabtay for their valuable feedback and assistance.
Appendix A Appendix
In this appendix, we describe the implementation details in Appendix A.1, qualitative examples in Appendix A.2, additional human annotations details in Appendix A.3 and additional experiments in Appendix A.4.
Model Epochs BS LoRA Rank LR Checkpoint
LLaVA-1.5-7B 3 8 64 0.0001 llava-hf/llava-1.5-7b-hf
TinyLLaVA 3 8 64 0.0001 tinyllava/TinyLLaVA-Gemma-SigLIP-2.4B
PaliGemma 15 8 128 0.0001 google/paligemma-3b-mix-224
Table 4: Training hyperparameters for the models. BS and LR refers to batch size and learning rate.
A.1 Implementation Details
KnowAda Implementation Details  For question generation, we use the gemini-1.5-flash-001 model, following the prompt in Figure 14. To generate answers, we set the temperature to 0.4 for all models, sampling 10 answers per question. These answers are evaluated using the gemini-1.5-flash-001 model, with prompts provided in Figure 15. For rewriting image descriptions, we employ gemini-1.5-pro, utilizing manually curated few-shot examples as illustrated in Figure 16. In all experiments, we use greedy sampling when generating outputs with Gemini.
Proposition Entailment Evaluation Implementation Details  We use gemini-1.5-flash-001 both for the proposition extraction, as well as the textual entailment task. We constrain the output to be in JSON format, following the proposition extraction prompt shown in Figure 17 and textual entailment prompt in Figure 18.
Training Details  We provide the training hyperparameters in Table 4, all other hyperparameters are set to default. LLaVA and PaliGemma were trained using LLaMA-Factory framework Zheng et al. (2024) and TinyLLaVA was trained using TinyLLaVA-Factory framework Jia et al. (2024). We note that this work fully complies with the licenses of all used scientific artifacts (e.g. DOCCI, LLaVA, PaliGemma, etc.). All use of scientific artifacts is consistent with their intended use. All models were train on a single NVIDIA A6000 GPU.
Figure 7: Distribution of number of unknown questions per caption for each ratio threshold of correct and incorrect answers from the vlm.
Figure 8: Distribution of number of questions per caption that are generated by the LLM based on the ground truth image description during the first stage of KnowAda.
Figure 9: Distribution of contradicted propositions’ relative location in generated captions across different models, tested on the DOCCI test set and measured using our proposition evaluation framework. Contradictions increase progressively as the captions become longer.
A.2 Qualitative Examples
In Figure 11, we show example of our evaluation pipeline, including the original description, the generated description and the labels produce by the NLI model.
Figure 12 and Figure13 shows outputs generated by LLaVA-1.5-7B, trained on KnowAda-adapted captions using various thresholds. As the threshold $T$ decreases, the model’s tendency to hallucinate decreases. However, this comes at the cost of reduced detail coverage. The threshold $T$ allows for fine-grained control over this trade-off between hallucination and detail retention.
A.3 Additional Annotation Details
To recruit high-quality annotators, we required them to have a HIT rate greater than 97% and at least 5,000 approved HITs, without imposing any geographical constraints. Additionally, we designed a qualification test consisting of 10 questions of varying difficulty, which we manually evaluated. A total of 14 annotators who passed the test with fewer than 2 mistakes were selected to perform the annotations for all experiments.
We present an example from the Amazon Mechanical Turk user interface, as shown to the annotators in Figure 10. Alongside the generated task guidelines, we provided supplementary slides with a detailed explanation of the visual entailment task, including 10 manually annotated examples that feature different labels and varying levels of difficulty for the annotators to review.
Annotators were paid 0.10\$ per annotation, with an average hourly wage of 10\$. The annotation process was approved by the institution’s ethics committee.
A.4 Additional Experiments
Relative Location of Contradicted Propositions. In Figure 9, we illustrate the distributions of the relative locations of propositions categorized as contradicted with respect to the generated caption. Consistent with prior research McKenna et al. (2023) We observe that errors tend to increase with the distance from the beginning of the text: as more text is written, the rate of contradictions increases.
DNLI Evaluation with a Smaller Model To justify the use of Gemini in the question difficulty evaluation stage, we compared its performance against Gemma-2B by using both models as judges on 1,000 question-answer pairs from the DOCCI test set. The results showed a 75% agreement between the two models. To further analyze discrepancies, we manually inspected 20 random samples where the judgments differed, finding that in 18 out of 20 cases, Gemini provided the accurate assessment. These findings highlight a trade-off between model size and judgment accuracy. While Gemma-2B is a more lightweight alternative, we prioritized the superior accuracy and consistency offered by Gemini for our evaluations.
Additional KnowAda Statistics  In Figure 8, we show the distribution of the number of questions generated per caption for the DOCCI training set. The data indicates that the distribution is centered around approximately 13 visual questions per caption.
In Figure 7, we demonstrate how the number of unknown questions changes when adjusting the threshold $T$. As T increases, more questions are classified as unknown, shifting the distribution.
Figure 10: Example of annotation interface in Amazon Mechanical Turk
Figure 11: Qualitative Example of Our Proposition-Based Evaluation Metric. Given a generated description by a VLM (middle), we decompose it into atomic propositions (top, center part). Then, we classify each proposition as either ’Entailed’, ’Contradicted’ or ’Neutral’ (top, right part), conditioned on the ground-truth image description (bottom). Finally, we calculate the consistency and contradiction based on the number of entailed and contradicted propositions.
Figure 12: Qualitative Example of KnowAda’s Threshold $T$ effect. We apply KnowAda’s rephrasing to the original caption (bottom) using different TH values.
Figure 13: Additional Qualitative Example of KnowAda’s Threshold $T$ effect.
Figure 14: Question Generation Prompt.
Figure 15: Evaluate VLM Answer Prompt.
Figure 16: Rewriting Prompt.
Figure 17: Proposition Extraction Prompt.
Figure 18: Proposition Judgement Prompt.